PPP Company Limited provides payday loan for an administration fee of $8 per loan, plus 5% of  the principal borrowed, plus 50% effective annual rate of interest. RRR Company offers payday loan for a flat fee of $10 per contract, plus 59% effective annual rate of interest. Anne Chang wants to borrow $500 for 15 days.

Required for each company: (a) what is the cost in dollars of the loan? (8 marks) (b) what is the actual EAR of the loan (2 marks)?
Answer #1

Amount borrowed is $500

Number of days = 15 days

a.> PPP company, cost of Dollars of the loan:

Admin. fee = $ 8 per loan

Additional fee = 5% of loan

= $ 500*5% = $ 25

Effective Interest rate = 50%

Interest on borrowed amount = $500*50%*(15/365)

= $ 10.27

Total Cost in dollars of loan = $ 8 + $ 25 +$ 10.27

= $43.27

Now, RRR company, cost of Dollars of the loan:

Fee per contract = $ 10 per loan

Additional fee = $0

Effective Interest rate = 59%

Interest on borrowed amount = $500*59%*(15/365)

= $ 12.12

Total Cost in dollars of loan = $ 10+ $ 12.12

= $22.12

b.> Actual EAR of the loan for each company:

For PPP Company

Total Cost = $ 43.27

Total Cost = amount Borrowed*EAR*(15/365)

43.27 = 500*EAR%*(15/365)

EAR = 210.58%

For RRR Company

Total Cost = $ 22.12

Total Cost = amount Borrowed*EAR*(15/365)

22.12 = 500*EAR%*(15/365)

EAR = 107.65%